solve the system of equations using a matrix. separate the x and the y values with a comma. 2x+y=-22, -3x+8y=-24

Answer by Alan3354(69443)

2x+y=-22
-3x+8y=-24
------------
|+2 +1 22|
|-3 +8 24|
----------
Det = 2*8 - (-3)*1
Det = 19
--------
x*Det = 1*24 - 8*22 = -152
x = -8
-----------
-y*Det = 2*24 + 66 = 114
y = -6